# Asynchronous Predictive Counterfactual Regret Minimization+ Algorithm in Solving Extensive-Form Games

### v1

@@ Appendix D Additional Experiments
Additionally, we find that both $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ and $\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$ decrease for our algorithm compared to PCFR+. We speculate that this is due to our algorithms’ higher stability compared to PCFR+. Specifically, for the reduction of $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$, when updating using the prediction, our algorithm utilizes smaller learning rates, resulting in smaller gaps between strategies indicated by different explicit accumulated regrets. Intuitively, it leads to smaller value of $\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$. For the reduction of $\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$, we attribute it to that as the learning rate decreases, the gap between the strategies represented by $\bm{R}^{t}_{I}$ and $\hat{\bm{R}^{t}_{I}}$ also diminishes, which implies the $\bm{r}^{t}_{I}$ and instantaneous counterfactual regret derived from the strategy represented by $\bm{R}^{t}_{I}$ becomes closer. This leads to more stable updates from $\bm{R}^{t}_{I}$ to $\bm{R}^{t+1}_{I}$.
@@
Furthermore, we observe that the improvement of our algorithms over PCFR+ is correlated with the rate at which the inaccuracy $\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ between the predicted and observed instantaneous counterfactual regrets decreases in PCFR+. When considering the game size, i.e., the number of the infosets (as shown in Table 2), we find that this inaccuracy decreases particularly rapidly in Battleship (3) and Goofspiel (4). Specifically, the value of the accumulated inaccuracy $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ stops increasing after approximately 10 iterations in only Kuhn Poker, Battleship (3), and Goofspiel (4), with Kuhn Poker having significantly fewer infosets than Battleship (3) and Goofspiel (4). According to the experimental results in Figure 3, these games are also the games where our algorithms do not outperform PCFR+.
@@
Dynamics of $\sum_{t=1}^{T}(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2})$. The dynamics of $\sum_{t=1}^{T}(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2})$ for our algorithms and PCFR+, which is related to the upper bound of the counterfactual regret presented in Theorem 4.1, are shown in Figures 10 and 11. These figures show the cumulative values across all infosets. We observe that the values of our algorithms, APCFR+ and SAPCFR+, are lower than those of PCFR+. This indicates that our algorithms achieve a lower upper bound on the counterfactual regret. Notably, we find that for most cases, the values of this term are smaller in APCFR+ than in SAPCFR+, with exceptions occurring only in Goofspiel (4) and Subgame 4. A lower upper bound on the counterfactual regret implies a faster convergence rate, which is consistent with the experimental results presented in Figures 3 and 4.
@@
Dynamics of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{{\bm{r}}^{t-1}_{I}}{{1+\alpha^{t}_{I}}}\|^{2}_{2}$. The dynamics of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{{\bm{r}}^{t-1}_{I}}{{1+\alpha^{t}_{I}}}\|^{2}_{2}$ for both our algorithms and PCFR+, which is related to the upper bound on counterfactual regret presented in Theorem B.1, is illustrated in Figures 12 and 13. These figures present the cumulative values across all infosets. For PCFR+, the expressions $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{{\bm{r}}^{t-1}_{I}}{{1+\alpha^{t}_{I}}}\|^{2}_{2}$ and $\sum_{t=1}^{T}\left(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\right)$ are equivalent, as $\alpha^{t}_{I}=0$ for PCFR+. By synthesizing the results from Figures 10, 11, 12, and 13, we observe that for our algorithms, the value of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{{\bm{r}}^{t-1}_{I}}{{1+\alpha^{t}_{I}}}\|^{2}_{2}$ significantly exceeds that of $\sum_{t=1}^{T}\left(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\right)$ as the number of iterations increases. Specifically, we observe that the value of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}$ consistently increases over time. In contrast, the value of $\sum^{T}_{t=1}\left(\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}\right)$ tends to stabilize, exhibiting a flattening trend. We hypothesize that this phenomenon arises because, even when ${\bm{r}}^{t}_{I}$ and $\bm{r}^{t-1}_{I}$ are close, the value of $\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}$ remains significantly large. Consequently, we employ Theorem 4.1 in the main text instead of Theorem B.1.

### v2

@@ Appendix D Additional Experiments
Additionally, we find that both $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ and $\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$ decrease for our algorithms (APCFR+ and SAPCFR+) compared to PCFR+. We speculate that this is due to our algorithms’ higher stability compared to PCFR+. Specifically, for the reduction of $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$, when updating using the prediction, our algorithm utilizes smaller learning rates, resulting in smaller gaps between strategies indicated by different explicit accumulated regrets. Intuitively, it leads to smaller value of $\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$. For the reduction of $\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$, we attribute it to that as the learning rate decreases, the gap between the strategies represented by $\bm{R}^{t}_{I}$ and $\hat{\bm{R}^{t}_{I}}$ also diminishes, which implies the $\bm{r}^{t}_{I}$ and instantaneous counterfactual regret derived from the strategy represented by $\bm{R}^{t}_{I}$ becomes closer. This leads to more stable updates from $\bm{R}^{t}_{I}$ to $\bm{R}^{t+1}_{I}$.
@@
Furthermore, we observe that the improvement of our algorithms over PCFR+ is correlated with the rate at which the inaccuracy $\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ between the predicted and observed instantaneous counterfactual regrets decreases in PCFR+. When considering the sizes of the games (Table 3), this inaccuracy decreases particularly rapidly in Battleship (3,2,3) and Goofspiel (4). Specifically, the value of the accumulated inaccuracy $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ stops increasing after approximately 10 iterations in only Kuhn Poker, Battleship (3,2,3), and Goofspiel (4), with Kuhn Poker is significantly smaller than Battleship (3,2,3) and Goofspiel (4). According to the experimental results in Figure 3, these games are also the games where our algorithms outperform similar with PCFR+.
@@
Dynamics of $\sum_{t=1}^{T}(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2})$. The dynamics of $\sum_{t=1}^{T}(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2})$ for our algorithms and PCFR+, which is related to the upper bound of the counterfactual regret presented in Theorem 4.1, are shown in Figures 8 and 9. These figures show the cumulative values across all infosets. We observe that the values of our algorithms, APCFR+ and SAPCFR+, are lower than those of PCFR+. This indicates that our algorithms achieve a lower upper bound on the counterfactual regret. Notably, we find that for most cases, the values of this term are smaller in APCFR+ than in SAPCFR+, with exceptions occurring only in Goofspiel (4) and Subgame 4. A lower upper bound on the counterfactual regret implies a faster convergence rate, which is consistent with the experimental results presented in Figure 3 and Table 1.
@@
Dynamics of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{{\bm{r}}^{t-1}_{I}}{{1+\alpha^{t}_{I}}}\|^{2}_{2}$. The dynamics of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{{\bm{r}}^{t-1}_{I}}{{1+\alpha^{t}_{I}}}\|^{2}_{2}$ for both our algorithms and PCFR+, which is related to the upper bound on counterfactual regret presented in Theorem B.1, is illustrated in Figures 10 and 11. These figures present the cumulative values across all infosets. For PCFR+, the expressions $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{{\bm{r}}^{t-1}_{I}}{{1+\alpha^{t}_{I}}}\|^{2}_{2}$ and $\sum_{t=1}^{T}\left(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\right)$ are equivalent, as $\alpha^{t}_{I}=0$ for PCFR+. By synthesizing the results from Figures 8, 9, 10, and 11, we observe that for our algorithms, the value of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{{\bm{r}}^{t-1}_{I}}{{1+\alpha^{t}_{I}}}\|^{2}_{2}$ significantly exceeds that of $\sum_{t=1}^{T}\left(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\right)$ as the number of iterations increases. Specifically, we observe that the value of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}$ consistently increases over time. In contrast, the value of $\sum^{T}_{t=1}\left(\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}\right)$ tends to stabilize, exhibiting a flattening trend. We hypothesize that this phenomenon arises because, even when ${\bm{r}}^{t}_{I}$ and $\bm{r}^{t-1}_{I}$ are close, the value of $\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}$ remains significantly large. Consequently, we employ Theorem 4.1 in the main text instead of Theorem B.1.
Empirical convergence rates of APCFR+ with an alternative learning approach for $\alpha^{t}$. In Eq (10), we proposed a learning approach for $\alpha^{t}$. Here, we evaluate an alternative learning approach for $\alpha^{t}$, specifically defined as $\alpha^{t}_{I}=\min\left(\sqrt{\frac{\max_{\tau\in[t-1]}\|\bm{r}^{\tau}_{I}-\bm{r}^{\tau-1}_{I}\|^{2}_{2}}{\max_{\tau\in[t-1]}\|\bm{R}^{\tau+1}_{I}-\bm{R}^{\tau}_{I}\|^{2}_{2}}},\ \alpha_{max}\right)$. The experimental results are presented in Figure 12 and Table 5, where the APCFR+ with new learning approach for $\alpha^{t}$ is denoted as “APCFR+ V2”. These results demonstrate that this alternative $\alpha^{t}$ learning approach generally achieves performance comparable to the method outlined the original APCFR+ in most games. We also observe that this alternative $\alpha^{t}$ learning approach slightly outperforms the original APCFR+ in large games such as Battleship (3,2,3), Battleship (4,3,2), Subgame (3), and Subgame (4). A potential direction for future work is to develop more efficient algorithms for solving large IIGs, building upon the new $\alpha^{t}$ learning approach. However, in Liar’s Dice (5), this alternative approach performs significantly worse than the method in Eq. (10).
Comparison with other classical CFR algorithms and the generalization of our approach. Now, we evaluate our algorithms with other classical CFR algorithms such as CFR (zinkevich2007regret), CFR+ (tammelin2014solving), and DCFR (brown2019solving). Initially, we conducted experiments using standard commonly used IIG benchmarks and HUNL Subgames. The experimental results are shown in Figure 13 and Table 6, respectively. We observed that APCFR+ and SAPCFR+ consistently outperformed CFR and CFR+ across all games. However, in poker games like Leduc Poker and HUNL Subgames, APCFR+ and SAPCFR+ did not surpass DCFR. Notably, our algorithms and DCFR are not mutually exclusive and can be combined effectively. Specifically, the core idea of our algorithm—the asymmetry of step sizes—can be integrated with that of DCFR, which discounts prior iterations when calculating accumulated regrets. Consequently, we apply the asymmetry mechanism for step sizes to enhance PDCFR+ (xu2024minimizing), a combination of DCFR and PCFR+, resulting in APDCFR+. We only make evaluation in HUNL Subgames as Leduc Poker is extremely smaller than HUNL Subgames. The experimental results are presented in Table 6. APDCFR+ significantly outperforms both PDCFR+, APCFR+, SAPCFR+, and DCFR (note that APCFR+ and SAPCFR+ outperform PDCFR+). We also compare APDCFR+ with DCFR+ (xu2024minimizing), an advanced variant of DCFR. The experimental results in Table 6 indicates that APDCFR+ also significantly outperforms DCFR+.
As mentioned earlier in our main text, we further evaluate the performance of DCFR, PCFR+, APCFR+, SAPCFR+, DCFR+, and APDCFR+ on Leduc Poker variants, which is larger than the vanilla Leduc Poker. Detailed experimental results and discussions can be found in the final paragraph of Section 5 in the main text.
Appendix E Limitations
The primary limitation of APCFR+ and SAPCFR+ is the theoretical convergence rate of $O(1/\sqrt{T})$. Notably, nearly all CFR algorithms exhibit a theoretical convergence rate of $O(1/\sqrt{T})$. Only a few CFR algorithms have a theoretical convergence rate exceeding $O(1/\sqrt{T})$, e.g., Reg-CFR and Clairvoyant CFR, which achieve $O(1/T^{\frac{3}{4}})$ and $O(1/T)$ theoretical convergence rates, respectively. Unfortunately, as shown in our experiments, the empirical convergence rate of these algorithms fails to surpass that of CFR algorithms, which have a theoretical convergence rate of $O(1/\sqrt{T})$, e.g., PCFR+. Interestingly, to achieve a faster empirical convergence rate, some CFR algorithms like DCFR, intentionally compromise their theoretical convergence rate. Specifically, the theoretical convergence rate of DCFR is 3 times slower than that of CFR+ (under the same number of iterations, the theoretical upper bound on the exploitability of DCFR is three times higher than that of CFR+), as shown in Theorem 3 of tammelin2015solving (CFR+) and Theorem 2 of brown2019solving (DCFR), while DCFR usually outperforms CFR+, as shown in our experiments.
@@ Appendix E Limitations
Figure 12: Empirical convergence rates of PCFR+, APCFR+, SAPCFR+, and APCFR+ V2 in standard commonly used IIG benchmarks.
PCFR+ APCFR+ SAPCFR+ APCFR+ V2
S3 1.44e-3 1.02e-3 (-29.2%) 9.44e-4 (-34.4%) 9.84e-4 (-31.7%)
S4 1.04e-3 7.53e-4 (-27.6%) 7.83e-4 (-24.7%) 7.38e-4 (-29.0%)
Table 5: The final exploitability for PCFR+, APCFR+, SAPCFR+, and APCFR+ V2 in HUNL Subgames. Values in red indicate percentages relative to PCFR+.
Figure 13: Empirical convergence rates of CFR, CFR+, DCFR APCFR+, and SAPCFR+ in standard commonly used IIG benchmarks.
CFR CFR+  DCFR  APCFR+ x SAPCFR+ x PDCFR+
S3 1.96e-2 1.16e-2 3.05e-4 1.02e-3 9.44e-4 1.08e-3
S4 2.68e-2 1.82e-2 2.18e-4 7.53e-4 7.83e-4 1.10e-3
DCFR+ APDCFR+
S3 2.23e-4 (-26.8%) 1.69e-4 (-44.6%)
S4 1.55e-4 (-28.9%) 1.27e-4 (-41.7%)
Table 6: The final exploitability for CFR, CFR+, and APDCFR+ in HUNL Subgames. Values in blue indicate percentages relative to DCFR. “S3” and “S4” denote Subgame 3 and Subgame 4, respectively.